# What Determines the Maximum Mass of AGN-assisted Black Hole Mergers?

### v1

@@ 3.1 Life of black hole binaries
We illustrate the interaction rate of binary or single BHs in Figure 8. For this analysis, we use SMBH mass $4\times 10^{6}M_{\odot}$, with the background disk parameters evaluated at $t=3$ Myr, adopting a Gaussian inclination distribution with $\beta_{v}=0.2$. The single BH mass is set to $10M_{\odot}$, while the BH binary mass two BHs is set to $10M_{\odot}+10M_{\odot}$. The single and binary BHs are assumed to move at relative velocities $v_{x,y,z}=10^{-3}c_{s}/\sqrt{3}$ (inside the disk components), $v_{x,y,z}=10^{-1}c_{s}/\sqrt{3}$ (inside the disk but outside the disk components) or $v_{x,y,z}=10^{-1}v_{\mathrm{kep}}/\sqrt{3}$ (outside the disk), where $c_{s}=v_{\mathrm{kep}}(h/r)$ represents the local sound speed of the AGN disk. The choice of outer-disk BH velocity indicates the inclination angle $i=0.058\,\mathrm{rad}=3.3^{\circ}$. The binary’s separation is taken as the Hill radius $r_{\mathrm{Hill}}(r,M=20M_{\odot})$, which is the most newly formed binaries’ separation, $10^{-2}r_{\mathrm{Hill}}$, and a fixed separation $10^{-6}$ pc. In Fig. 8, we account for the fraction of time the BHs spend within the AGN disk $p_{\mathrm{disk}}$ or within the disk components $p_{c}$. Additionally, we set $p_{\rm{uni}}=0.5$ in the shear velocity equation (Eq. 69) to mitigate randomness.
@@
In Figure 8 (a), the interaction rates for a BH in the AGN disk reveal that gas dynamical friction (green) generally dominates over other interactions. The accretion rate (orange) and migration rate (pink) are typically around 1000 times smaller than the gas dynamical friction rate. For BHs located outside the disk (dotted line in Figure 8 (a) and (c)), the interaction rates are lower due to the factor $p_{\mathrm{disk}}$, particularly for binary formation due to gas capture, which is more likely within the disk environment. Notably, gas dynamical friction does not operate effectively at certain velocities due to the feedback condition Eq. 52, leading to a disappearance in the green line at larger radii, observed in both panels (a) and (c).
@@
In Figure 8 (b), we plot the hardening rate for BH binaries within the disk, as gas hardening occurs only in the disk. When a binary forms at a separation of $s\sim r_{\mathrm{Hill}}$, gas dynamical friction plays a dominant role in hardening the system. As the separation decreases quickly to a smaller value $10^{-2}r_{\mathrm{Hill}}$, the effectiveness of dynamical friction diminishes, while interactions with the circum-binary disk become increasingly significant. These two processes continue to shrink the binary orbit until gravitational wave emission eventually takes over, driving a rapid merger. When the BH binary migrates to smaller radii, the Hill radius $r_{\mathrm{Hill}}$ decreases, resulting in a stronger gas dynamical friction rate. Consequently, the transition from gas-hardening dominance to gravitational-wave dominance occurs over a shorter timescale, effectively accelerating the hardening process (e.g. dash-dot line at small radii) . This progression aligns well with our previous understanding of binary mergers in AGN disks.
@@
In Figure 8 (c), binaries form at a very high rate within the AGN disk, primarily due to gas capture. The time fraction spent in the disk components, $p_{c}$, has only a minor effect on the interaction rate (solid and dashed lines) when the BH is inside the disk. This process provides an abundant supply of binaries within the AGN disk, facilitating the formation of higher-generation BH binaries from remnant single BHs.
@@
During the migration and hardening process, binary-single interactions with disk stars play a crucial role in shaping binary dynamics. As shown in Figure 8 (d), interactions with background spherical stars (green line) regularly soften newly formed binaries ($s\sim r_{\mathrm{Hill}}$) at large radii. However, as the binary separation decreases, the impact of these interactions diminishes significantly. The interaction rate with disk BHs (red line) and disk stars (blue line) remains relatively unaffected by binary separation and the relative velocity $v$. Within the AGN disk, binaries frequently undergo binary-single interactions with disk BHs, especially at smaller radii. These encounters rapidly harden the binary but can also eject it from the disk, effectively halting gas-driven hardening and slowing its inward migration toward the center.
@@
We can then construct a binary evolution pathway similar to that presented in Figure 5 of Tagawa et al. (2020b). In AGN disks, BH binaries can form at larger radii, migrate inward, while rapidly hardening through gas dynamical friction. As the binary separation decreases, gas dynamical friction becomes less effective, and circum-binary disk interactions take over as the dominant hardening mechanism. At very small separations, gravitational wave emission ultimately governs the final stage of hardening, leading to the binary’s merger into a single BH. Binaries at smaller radii typically transition more quickly from the gas dynamical friction hardening phase to the gravitational wave-driven hardening phase. However, during the inward migration, binaries may undergo multiple binary-single interactions with other disk objects, which can further harden the system but eject the binary from the AGN disk. Ejected binaries can often re-enter the disk relatively quickly, resuming their inward migration and continuing to harden through gas interactions. This dynamic interplay of migration, hardening, and interactions shapes the overall evolution and merger likelihood of BH binaries in AGN disks. Our results align well with those of Tagawa et al. (2020b).
@@ 3.2 Merger features
In the binary formation channel, Eqs. 94 and 104 introduce a double-counting effect. To correct for this, we assign a weighting factor to each merger, given by the inverse of the remnant BH’s generation $w=1/\rm{gen}$. The classification of BH generations is defined as follows: pre-existing binaries and binaries formed within the AGN disk are designated as the 0th generation. After a merger, their remnant single BHs become 1st generation BHs. Additionally, pre-existing single BHs and those originating from the AGN disk are also classified as 1st generation. When an $n$th generation BH forms a binary and subsequently merges, the remnant single BH is classified as $(n+1)$th generation.
We primarily consider two inclination (velocity) models: an isotropic distribution and an anisotropic Gaussian distribution with $\beta_{v}=0.2$. These two models behave very different in the number of BHs embedded in the disk due to the inclination difference, as shown in Fig. 3. Fig. 9 illustrates the radial and time distribution of mergers for both the Gaussian model (top two rows) and the isotropic model (bottom two rows), categorized by binary formation channel or remnant generation. We assume an AGN lifetime of 100 Myr; however, in our model, the AGN lifetime solely determines the simulation duration and does not affect the AGN disk structure or BH distribution properties. Consequently, the characteristics of mergers in a shorter AGN lifetime scenario can also be inferred from the figure. The averaged Merger rate and top 1% merger mass is plotted in Fig. 10.
It is evident that mergers originating from gas-captured binaries dominate the merger events, occurring at rates approximately $\mathcal{O}(10)$ times higher than those of pre-existing binary mergers, both with a slight decline over time. This decrease occurs as disk BHs gradually deplete, while mergers from the dynamical formation and star formation channels continue at a steady rate. Overall, the merger rate across all binary types remains roughly constant, aligning well with Fig. 7 in Tagawa et al. (2020b).
@@
The merger events are clustered at approximately $10^{-3}$ pc at early time, with their positions being closer to the SMBH in the isotropic model. These regions of frequent mergers generally correspond to disk BH number densities of $n\approx 10^{13}-10^{14}\;\rm{pc^{-3}}$. Such high number densities (calculated by Eq.120) arise from the extremely low scale height of the disk BH component, as discussed in Section 2.6.2. As time progresses, the location of the critical density shifts outward as shown in the number density map, causing the gathered merger positions to increase correspondingly in the density heatmap. A small bump is observed in the merger position figure due to the sudden slow down of migration speed at $r\sim 10^{-3}$ pc (a discontinuous migration trap), which is also visible in the density heatmap of merger time and mass.
Merger remnant generations can reach extremely high values, with a maximum of approximately 100 observed at late times. The highest-generation mergers closely follow their merger time and location, as these BHs require extended time in dense regions of the disk BHs to undergo successive mergers. Mergers of generation greater than 64 are concentrated in the region $10^{-3}-10^{-2}$ pc, where the high disk BH density facilitates frequent binary-single interactions, providing a continuous supply of binary seeds. Moreover, the high gas density in this region enables BHs to rapidly re-enter the disk, accelerating the hardening and merger process. These very high-generation BHs are typically form at late times ($t\gtrsim 80$ Myr), with their merger masses growing significantly through repeated mergers.
The maximum merger mass increases exponentially with AGN lifetime, as shown in the right panel of Fig. 10. In the density heatmap, we trace the merger history of the most massive mergers before 10, 40, 70, and 100 Myr. The migration speeds of these BHs progressively slow down due to increasing disk BH number densities over time. At $t\lesssim 40$ Myr, the maximum masses in the Gaussian and isotropic models differ slightly, but they converge at long AGN lifetimes, primarily due to the early-time differences in BH number density.
Figure 10: The averaged merger rate and highest 1% merger mass as a function AGN lifetime $t_{\mathrm{AGN}}$. The blue line represent the merger rate for a Gaussian distribution inclination, and the orange line is for an isotropic distribution inclination. The merger rate figure is the averaged merger rate of Fig. 9.
@@ 3.3 Merger rate and highest mass
Table 2: The parametric relevance $R$ (defined in Eq. 135) for pre-existing BHs (including the star forming ”pre-existing” binaries) during the simulation under isotropic and Gaussian distributions of inclination. The plot range of each parameter is listed in Table 1. The red cells represent the parameter plays an important role ($R>0.75$) and the blue cells means this parameter is relevant but less important ($0.750.4<R<$ 0.75) to the merger rate / highest mass for Gaussian / isotropic distribution. We estimate the relevance of $(1+\gamma_{\rho})$ instead of $\gamma_{\rho}$ to remove the singularity in Eq. 135.
Figure 11: The averaged merger rate as a function of different parameters listed in Table 1 (except for $t_{\mathrm{AGN}}$). The blue line represent the merger rate for a Gaussian distribution inclination, and the orange line is for an isotropic distribution inclination.
@@
We then estimate the average merger rate during the AGN disk lifetime (in $\mathrm{Myr}^{-1}\cdot\mathrm{AGN^{-1}}$) and average mass of the top 1% merger mass ($M_{\mathrm{BH}}=M_{1}+M_{2}$ in $M_{\odot}$) according to their merger generations across different parameter configurations listed in Table 1. The average merger rate is given by:
@@
while the average mass of the top 1% of mergers is computed as:
@@
where $A$ is the merger rate $\Gamma_{\mathrm{merger}}$ or highest mass $M_{\mathrm{top}}$, and $f$ denotes the parameters varied in the simulation. A parameter is considered significant if $R\gtrsim 1$ (colored in red in Table 2), while irrelevant parameters typically correspond to cases where $R\sim\mathcal{O}(0.1)$. Parameters with moderate relevance $R\approx 0.5$, are colored in blue in Table 2. The role of AGN life time is discussed in Sec 3.2, with the trends plotted in Fig 10.
@@ 3.3.1 Impact of gas interaction
Faster migration speeds (larger $f_{\mathrm{mig}}$) mean that BHs spend less time in low gas density regions, thereby accelerating their mergers. Additionally, BHs concentrated in high-density gas regions experience more frequent binary-single interactions, which enhance binary hardening but can also reduce migration speed by ejecting BHs. As illustrated in the upper 2 lines of Fig. 13, the merger time decreases and the merger position gets closer to the center as the migration speeds up.
Interestingly, $f_{\mathrm{mig}}$ appears to have little effect on the highest merger mass in the isotropic distribution. While increasing migration speed slightly increases the maximum merger mass in the Gaussian model, the isotropic model reduces the maximum mass slightly. As seen in Fig. 13, high-generation mergers of isotropic model occur at smaller radii, very close to $r_{\mathrm{in}}$, and exhibit little dependence on $f_{\mathrm{mig}}$ when $f_{\mathrm{mig}}>2$. Additionally, in the Gaussian model, no BHs are observed migrating into $r<r_{\mathrm{in}}$ region, whereas in the isotropic model, the BHs migrating into $r<r_{\mathrm{in}}$ increases as $f_{\mathrm{mig}}$ increases. This difference arises due to the larger disk BH number density in the Gaussian model, leading to more frequent binary single interactions and the increased potential for higher-generation mergers. Notably, under realistic conditions, $f_{\mathrm{mig}}$ varies with radial distance on a small scale but remains around $f_{\mathrm{mig}}\sim 2.0$, as discussed in Sec. 4.3.

### v2

@@ 3.1 Life of black hole binaries
We illustrate the interaction rate of binary or single BHs in Fig. 3. For this analysis, we use SMBH mass $4\times 10^{6}M_{\odot}$, with the background disk parameters evaluated at $t=3$ Myr, adopting a Gaussian inclination distribution with $\beta_{v}=0.2$. The single BH mass is set to $10M_{\odot}$, while the BH binary mass of two BHs is set to $10M_{\odot}+10M_{\odot}$. The single and binary BHs are assumed to move at relative velocities $v_{x,y,z}=10^{-3}c_{s}/\sqrt{3}$ (inside the disk components), $v_{x,y,z}=10^{-1}c_{s}/\sqrt{3}$ (inside the disk but outside the disk components) or $v_{x,y,z}=10^{-1}v_{\mathrm{kep}}/\sqrt{3}$ (outside the disk), where $c_{s}=v_{\mathrm{kep}}(h/r)$ represents the local sound speed of the AGN disk. The choice of outer-disk BH velocity indicates the inclination angle $i=0.058\,\mathrm{rad}=3.3^{\circ}$. The binary’s separation is taken as the Hill radius $r_{\mathrm{Hill}}(r,M=20M_{\odot})$, which is the most newly formed binaries’ separation, $10^{-2}r_{\mathrm{Hill}}$, and a fixed separation $10^{-6}$ pc. In Fig. 3, we account for the fraction of time the BHs spend within the AGN disk $p_{\mathrm{disk}}$ or within the disk components $p_{c}$. Additionally, we set $p_{\rm{uni}}=0.5$ in the shear velocity equation (Eq. E5) to mitigate randomness.
@@
In Fig. 3 (a), the interaction rates for a BH in the AGN disk reveal that gas dynamical friction (green) generally dominates over other interactions. The accretion rate (orange) and migration rate (pink) are typically around 1000 times smaller than the gas dynamical friction rate. For BHs located outside the disk (dotted line in Fig. 3 (a) and (c)), the interaction rates are lower due to the factor $p_{\mathrm{disk}}$, particularly for binary formation due to gas capture, which is more likely within the disk environment. Notably, gas dynamical friction does not operate effectively at certain velocities due to the feedback condition Eq. D17, leading to a disappearance in the green line at larger radii, observed in both panels (a) and (c).
@@
In Fig. 3 (b), we plot the hardening rate for BH binaries within the disk, as gas hardening occurs only in the disk. When a binary forms at a separation of $s\sim r_{\mathrm{Hill}}$, gas dynamical friction plays a dominant role in hardening the system. As the separation decreases quickly to a smaller value $10^{-2}r_{\mathrm{Hill}}$, the effectiveness of dynamical friction diminishes, while interactions with the circum-binary disk become increasingly significant. These two processes continue to shrink the binary orbit until gravitational wave emission eventually takes over, driving a rapid merger. When the BH binary migrates to smaller radii, the Hill radius $r_{\mathrm{Hill}}$ decreases, resulting in a stronger gas dynamical friction rate. Consequently, the transition from gas-hardening dominance to gravitational-wave dominance occurs over a shorter timescale, effectively accelerating the hardening process (e.g. dash-dot line at small radii) . This progression aligns well with our previous understanding of binary mergers in AGN disks.
@@
In Fig. 3 (c), binaries form at a very high rate within the AGN disk, primarily due to gas capture. The time fraction spent in the disk components, $p_{c}$, has only a minor effect on the interaction rate (solid and dashed lines) when the BH is inside the disk. This process provides an abundant supply of binaries within the AGN disk, facilitating the formation of higher-generation BH binaries from remnant single BHs.
@@
During the migration and hardening process, binary-single interactions with disk stars play a crucial role in shaping binary dynamics. As shown in Fig. 3 (d), interactions with background spherical stars (green line) regularly soften newly formed binaries ($s\sim r_{\mathrm{Hill}}$) at large radii. However, as the binary separation decreases, the impact of these interactions diminishes significantly. The interaction rate with disk BHs (red line) and disk stars (blue line) remains relatively unaffected by binary separation and the relative velocity $v$. Within the AGN disk, binaries frequently undergo binary-single interactions with disk BHs, especially at smaller radii. These encounters rapidly harden the binary but can also eject it from the disk, effectively halting gas-driven hardening and slowing its inward migration toward the center.
@@
We can then construct a binary evolution pathway similar to that presented in Fig. 5 of Tagawa et al. (2020b). In AGN disks, BH binaries can form at larger radii, migrate inward, while rapidly hardening through gas dynamical friction. As the binary separation decreases, gas dynamical friction becomes less effective, and circum-binary disk interactions take over as the dominant hardening mechanism. At very small separations, gravitational wave emission ultimately governs the final stage of hardening, leading to the binary’s merger into a single BH. Binaries at smaller radii typically transition more quickly from the gas dynamical friction hardening phase to the gravitational wave-driven hardening phase. However, during the inward migration, binaries may undergo multiple binary-single interactions with other disk objects, which can further harden the system but eject the binary from the AGN disk. Ejected binaries can often reenter the disk relatively quickly, resuming their inward migration and continuing to harden through gas interactions. This dynamic interplay of migration, hardening, and interactions shapes the overall evolution and merger likelihood of BH binaries in AGN disks. Our results align well with those of Tagawa et al. (2020b).
@@ 3.2 Merger features
In the binary formation channel, Eqs. E30 and E40 introduce a double-counting effect. To correct for this, we assign a weighting factor to each merger, given by the inverse of the remnant BH’s generation $w=1/\rm{gen}$. The classification of BH generations is defined as follows: preexisting binaries and binaries formed within the AGN disk are designated as the zeroth generation. After a merger, their remnant single BHs become first generation BHs. Additionally, preexisting single BHs and those originating from the AGN disk are also classified as first generation. When an $n$th generation BH forms a binary and subsequently merges, the remnant single BH is classified as $(n+1)$th generation.
We primarily consider two inclination (velocity) models: an isotropic distribution and an anisotropic Gaussian distribution with $\beta_{v}=0.2$. These two models behave very differently in the number of BHs embedded in the disk due to the inclination difference, as shown in Fig. 12. Fig. 4 illustrates the radial and time distribution of mergers for both the Gaussian model (top two rows) and the isotropic model (bottom two rows), categorized by binary formation channel or remnant generation. We assume an AGN lifetime of 100 Myr; however, in our model, the AGN lifetime solely determines the simulation duration and does not affect the AGN disk structure or BH distribution properties. Consequently, the characteristics of mergers in a shorter AGN lifetime scenario can also be inferred from the figure. The averaged merger rate and top 1% merger mass is plotted in Fig. 5.
It is evident that mergers originating from gas-captured binaries dominate the merger events, occurring at rates approximately $\mathcal{O}(10)$ times higher than those of preexisting binary mergers, both with a slight decline over time. This decrease occurs as disk BHs gradually deplete, while mergers from the dynamical formation and star formation channels continue at a steady rate. Overall, the merger rate across all binary types remains roughly constant, aligning well with Fig. 7 in Tagawa et al. (2020b).
@@
The merger events are clustered at approximately $10^{-3}$ pc at early time, with their positions being closer to the SMBH in the isotropic model. These regions of frequent mergers generally correspond to disk BH number densities of $n\approx 10^{13}-10^{14}\;\rm{pc^{-3}}$. Such high number densities (calculated by Eq. F13) arise from the extremely low scale height of the disk BH component, as discussed in Appendix F.2. As time progresses, the location of the critical density shifts outward as shown in the number density map, causing the gathered merger positions to increase correspondingly in the density heatmap. A small bump is observed in the merger position figure due to the sudden slow down of migration speed at $r\sim 10^{-3}$ pc (a discontinuous migration trap), which is also visible in the density heatmap of merger time and mass.
Merger remnant generations can reach extremely high values, with a maximum of approximately 100 observed at late times. The highest-generation mergers closely follow their merger time and location, as these BHs require extended time in dense regions of the disk BHs to undergo successive mergers. Mergers of generation greater than 64 are concentrated in the region $10^{-3}-10^{-2}$ pc, where the high disk BH density facilitates frequent binary-single interactions, providing a continuous supply of binary seeds. Moreover, the high gas density in this region enables BHs to rapidly reenter the disk, accelerating the hardening and merger process. These very high-generation BHs are typically formed at late times ($t\gtrsim 80$ Myr), with their merger masses growing significantly through repeated mergers.
Figure 5: The averaged merger rate and highest 1% merger mass as a function AGN lifetime $t_{\mathrm{AGN}}$. The blue line represents the merger rate for a Gaussian distribution inclination, and the orange line is for an isotropic distribution inclination. The merger rate figure is the averaged merger rate of Fig. 4.
@@ 3.3 Merger rate and highest mass
Table 2: The parametric relevance $R$ (defined in Eq. 3) for preexisting BHs (including the star forming ”preexisting” binaries) during the simulation under isotropic and Gaussian distributions of inclination. The plot range of each parameter is listed in Table 1. $R>0.75$ indicates that the parameter plays a dominant role, $0.4<R<0.75$ suggests that the parameter is relevant but less influential in determining the merger rate or maximum BH mass for the Gaussian and isotropic distributions. We estimate the relevance of $(1+\gamma_{\rho})$ instead of $\gamma_{\rho}$ to remove the singularity in Eq. 3.
Figure 6: The averaged merger rate as a function of different parameters listed in Table 1 (except for $t_{\mathrm{AGN}}$). The blue line represents the merger rate for a Gaussian distribution inclination, and the orange line is for an isotropic distribution inclination.
@@
We then estimate the average merger rate during the AGN disk lifetime (in $\mathrm{Myr}^{-1}\cdot\mathrm{AGN^{-1}}$) and average mass of the top 1% merger mass ($M_{\mathrm{BH}}=M_{1}+M_{2}$ in $M_{\odot}$) according to their merger generations across different parameter configurations listed in Table 1. The average merger rate is given by
@@
while the average mass of the top 1% of mergers is computed as
@@
where $A$ is the merger rate $\Gamma_{\mathrm{merger}}$ or highest mass $M_{\mathrm{top}}$, and $f$ denotes the parameters varied in the simulation. A parameter is considered significant if $R\gtrsim 1$, whereas irrelevant parameters generally correspond to $R\sim\mathcal{O}(0.1)$. Parameters with moderate relevance $R\approx 0.5$. The corresponding results are listed in Table 2.
3.3.1 Impact of AGN lifetime
The AGN lifetime $t_{\mathrm{AGN}}$ is a unique disk parameter in that it only extends the simulation duration without modify any interaction strength or disk properties directly. The maximum merger mass increases exponentially with AGN lifetime, without affecting the merger rate a lot, as shown in the Fig. 5. In the density heatmap of Fig. 4, we trace the merger history of the most massive mergers before 10, 40, 70, and 100 Myr. The merger positions of 1st generation and the inward migration speeds of these BHs progressively slow down due to increasing disk BH number densities over time. At early times $t\lesssim 40$ Myr, the maximum BH masses differ slightly between the Gaussian and isotropic models. However, these differences diminish at longer AGN lifetimes, primarily reflecting disparities in BH number density during the initial stages of disk evolution.
Directly determining the ages of AGNs remains observationally challenging. Nevertheless, both theoretical frameworks and indirect observational approaches have been developed to estimate AGN lifetimes across cosmic time (Enoki et al., 2014; Schawinski et al., 2015; Adelberger & Steidel, 2005; Miller et al., 2003). Given the strong sensitivity of the maximum BH merger mass to AGN lifetime observed in our simulations, this relationship may offer a novel method for inferring the current age of AGN disks—measured from the onset of disk formation—based on the most massive binary mergers. This approach provides a complementary alternative to traditional optical diagnostics.
@@ 3.3.2 Impact of gas interaction
Faster migration speeds (larger $f_{\mathrm{mig}}$) mean that BHs spend less time in low gas density regions, thereby accelerating their mergers. Additionally, BHs concentrated in high-density gas regions experience more frequent binary-single interactions, which enhance binary hardening but can also reduce migration speed by ejecting BHs. As illustrated in the upper two lines of Fig. 8, the merger time decreases and the merger position gets closer to the center as the migration speeds up.
Interestingly, $f_{\mathrm{mig}}$ appears to have little effect on the highest merger mass in the isotropic distribution. While increasing migration speed slightly increases the maximum merger mass in the Gaussian model, the isotropic model reduces the maximum mass slightly. As seen in Fig. 8, high-generation mergers of isotropic model occur at smaller radii, very close to $r_{\mathrm{in}}$, and exhibit little dependence on $f_{\mathrm{mig}}$ when $f_{\mathrm{mig}}>2$. Additionally, in the Gaussian model, no BHs are observed migrating into the $r<r_{\mathrm{in}}$ region, whereas in the isotropic model, the BHs migrating into $r<r_{\mathrm{in}}$ increases as $f_{\mathrm{mig}}$ increases. This difference arises due to the larger disk BH number density in the Gaussian model, leading to more frequent binary single interactions and the increased potential for higher-generation mergers. Notably, under realistic conditions, $f_{\mathrm{mig}}$ varies with radial distance on a small scale but remains around $f_{\mathrm{mig}}\sim 2.0$, as discussed in Sec. 4.3.